OpenStudy (anonymous):

Can anybody please tell me what is the basic formula for Bulk Modulus, Thanks.

geerky42 (geerky42):

\[\Huge B = -V\cdot\dfrac{\Delta P}{\Delta V}\] Where \(V\) is volume, \(\Delta P\) is change in pressure, and \(\Delta V\) is change in volume. Equivalently \(B = -\rho\cdot\dfrac{d P}{d \rho}\) where \(\rho\) is density. Hope that's what you need.

OpenStudy (anonymous):

Thanks!

geerky42 (geerky42):

No problem